OpenStudy (anonymous):

Fine the inverse of the function. f(x)=square x-9

OpenStudy (anonymous):

\[y = x^2 - 9\] To find inverse, switch x & y :\[x = y^2 - 9\]\[y = \pm \sqrt{x-9}\]

OpenStudy (anonymous):

how would i write it as f^-1= for x

OpenStudy (anonymous):

\[f^{-1} (x) = \pm \sqrt{x-9}\]